OpenStudy (anonymous):

find the exact value of sin (-120)

OpenStudy (anonymous):

sin(-a) = -sin(a) sin(x) = sin(180-x) Gogogogog

OpenStudy (anonymous):

I don't understand that? is the answer -sin 60

OpenStudy (anonymous):

I don't understand that? is the answer -sin 60

OpenStudy (anonymous):

yes, sorry

OpenStudy (anonymous):

knowak stop player hating ¬_¬

OpenStudy (anonymous):

but, of course, you need to know the value of sin 60... \[-\sin 60 = -\frac{\sqrt{3}}{2} \] I think

OpenStudy (anonymous):

dont you have to add 360 to make the number positive first?